OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−4, 0) and (−5, 0) (−4, 0) and (5, 0) (4, 0) and (−5, 0) (4, 0) and (5, 0)

OpenStudy (ujjwal):

Put y=0. The value then obtained from the equation will give you x intercepts.

OpenStudy (anonymous):

huh? you just lost meh

OpenStudy (ujjwal):

put y=0 in given equation!

OpenStudy (anonymous):

-_-

OpenStudy (ujjwal):

And what does that mean?

OpenStudy (anonymous):

oh nothing

OpenStudy (anonymous):

\[ \large y=f(x)=x^2-x-20 \] Does this look more like a form you're comfortable with?, then do what @ujjwal suggested.

OpenStudy (anonymous):

a little better

OpenStudy (anonymous):

If you now apply what @ujjwal told you, because x intersects means that y has to be zero, these are the roots of the equation, the points where it crosses the x-axis, therefore y=0, you will get: \[ \Large x^2-x-20=0\] And this might even look a little bit better then the last one.